OpenStudy (anonymous):

What is the slope of the line 2x - 5y = 9?

OpenStudy (bahrom7893):

get this into y = form...

OpenStudy (bahrom7893):

2x-5y=9 -5y=9-2x -5y = -2x+9 y = (2/5)x - 9/5 So slope is M = 2/5